Name the x and y intercept for 4x-6y = -24
Found 2 solutions by jim_thompson5910, PRMath:
Answer by jim_thompson5910(13794)


x-intercept


To find the x-intercept, plug in and solve for x


Start with the given equation.


Plug in .


Multiply and 0 to get 0.


Simplify.


Divide both sides by to isolate .


Reduce.


So the x-intercept is .


------------------------------------------


y-intercept


To find the y-intercept, plug in and solve for y


Start with the given equation.


Plug in .


Multiply and 0 to get 0.


Simplify.


Divide both sides by to isolate .


Reduce.


So the y-intercept is .


Notice that if we graph the equation , we can see that the graph has the x-intercept and the y-intercept

Graph of

Answer by PRMath(57)

Name the x and y intercept for 4x-6y = -24

When you want the "y" intercept, solve for "y" by making the "x" = 0
When you want the "x" intercept, solve for "x" by making the "y" = 0.

Let's solve for the y intercept first:

4x - 6y = -24 (original equation)
4(0) - 6y = -24 (solve for "y" by making "x" =0)
0 - 6y = -24 (divide both sides by -6)
y = 4


Therefore, your "y" intercept is: (0, 4)

Now for the x intercept:

4x - 6y = -24 (original equation)
4x - 6(0) = -24 (solve for "x" by making "y" = 0)
4x = -24 (divide both sides by 4)
x = -6
Therefore, your "x" intercept is: (-6, 0)
Hope this helped you. :-)